OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#9051 — Paris Nexus
Incident Report for Web Cloud
Resolved
The peer-link between th2-11-n5 and gsw-11-n5 is down.
Our customers will go through th2-11-n5.

Update(s):

Date: 2013-08-01 04:08:07 UTC
The new link ADVA is working.

Date: 2013-08-01 04:07:14 UTC
peer-link and up-link were all switched on active equipement.

Date: 2013-08-01 04:06:38 UTC
peer-links are OK,now we are making the uplinks

Date: 2013-08-01 04:05:59 UTC
We mounted the 1st peer-link on nexus pair-12.

Date: 2013-07-31 23:19:30 UTC
We will start now, we will cut the secondary lines to prepare the switch over.

Date: 2013-07-30 23:57:39 UTC
We are now facing repetition issues on several optics DWDM equipping our DSL switches in Paris and used to connect our Globalswitch POPs and TH2. This has a several times impact on the traffic (many cutoffs from few seconds to several minutes depending on customers).

We have therefore decided to change the techno and move the links GSW <> TH2 in DWDM active ADVA on a first half the concerned links. This has to allow the stability of the infrastructure and avoid new cutoffs. A team will move tomorrow morning to Paris in order to prepare a new link. The switching will be done the following night on 31/07 to 01/08 from midnight.

Date: 2013-07-30 16:35:11 UTC
Due to the optical problem, we lost securisation between
th2-12-n5 and gsw-12-n5.
The customers continue to function on th2-12-n5.

Date: 2013-07-30 06:25:33 UTC
Everything is OK following the reboot.

Date: 2013-07-30 06:25:12 UTC
Now we have a problem on pair-10 gsw-10 and th2-10 , we are reloading th2-10.

Date: 2013-07-30 04:34:16 UTC
gsw-12-n5 is up now, we will keep monitoring.

Date: 2013-07-30 04:33:54 UTC
gsw-12-n5 is not up following the reboot, we will send a technician to gsw tomorrow morning,currently, our customers will keep going through th2-12.

Date: 2013-07-30 04:31:57 UTC
We will reboot gsw-12-n5.

Date: 2013-07-30 04:31:44 UTC
We have the same problem between th2-12-n5 and gsw-12-n5.

Date: 2013-07-30 00:14:27 UTC
We will reboot th2-11-n5,traffic loss expected.
Posted Jul 30, 2013 - 00:13 UTC